Im trying to help my son with algebra homework i dont remember how to do it the problem is y-3-9=30
Answer by mystic1306(20)

Get the y by itself by adding 3 and 9 to both sides.
y-3+3-9+9=30+3+9
as you see on the left -3+3 and -9+9 cancels out, solve for y
y=30+3+9
y=42